# U-PINet: Physics-Informed Hierarchical Learning for Accurate and Fast 3D RCS Prediction

### v2

@@ III-A Near-Field Graph Encoder Module
As shown in Fig. 4(a), the physics-guided edge kernel $\mathcal{G}(\cdot)$ is realized as an attention module that takes the geometric inputs and produces a near-field edge weight $w_{ij}$ for each neighboring pair $(i,j)$.
For each point $\boldsymbol{p}_{i}$, we estimate its local sampling density by Gaussian kernel [43] density estimation over $\mathcal{N}(i)$:
@@
where $N=|\mathcal{N}(i)|$ and $h>0$ is the kernel bandwidth with $\mathcal{K}(\boldsymbol{x})=\exp(-\|\boldsymbol{x}\|^{2}/2)$. The inverse density is
@@
where $F_{d}(\cdot)$ is a two-layer feed-forward network (FFN).
Local geometric anisotropy is encoded by the difference of unit normals
$\Delta\boldsymbol{n}_{ij}=\boldsymbol{n}_{i}-\boldsymbol{n}_{j},$ (29)
which yields a normal-based attention
@@
where $F_{n}(\cdot)$ is another two-layer FFN and $\{\gamma_{ij}\}_{j\in\mathcal{N}(i)}$ forms the normal-vector attention branch.
Finally, the near-field edge weight is obtained by combining the two branches:
$w_{ij}=\mathrm{MLP}\!\left(\beta_{ij}\;\oplus\;\gamma_{ij}\right),$ (31)
where $\oplus$ denotes feature concatenation and $\mathrm{MLP}(\cdot)$ is a multi-layer perceptron. This physics-guided edge attention provides a learnable realization of $\mathcal{G}(\cdot)$ in the near-field formulation, linking local density variation and geometric anisotropy to effective self- and mutual-impedance behavior.
@@
The network implementation of the near-field operator in Section II directly follows the iterative message-passing scheme in (8)-(10). Leveraging the graph topology and physics-embedded edge weights $\boldsymbol{w}_{ij}$ established in the construction phase, each iteration realizes the kernel $h_{t}(\boldsymbol{w}_{ij},\boldsymbol{J}^{(t)}(\boldsymbol{p}_{j}))$ using a lightweight stack of attention- and convolution-based graph layers as shown in Fig. 4(b).
Let $\mathbf{H}^{(0)}=\{\boldsymbol{J}^{(t)}(\boldsymbol{p}_{i})\}$ denote the initial current features. In particular, the module integrates Graph Attention Networks (GAT) [44] and Graph Convolutional Networks (GCN) [45]. One update cycle is implemented as
$\mathbf{H}^{(1)}$ $=\mathrm{ELU}\!\left(\mathrm{BN}\!\left(\mathrm{GAT}\big(\mathbf{H}^{(0)},\boldsymbol{w}\big)\right)\right),$ (32)
$\mathbf{H}^{(2)}$ $=\mathrm{ELU}\!\left(\mathrm{BN}\!\left(\mathrm{GCN}\big(\mathbf{H}^{(1)},\boldsymbol{w}\big)\right)\right),$
$\mathbf{H}^{(3)}$ $=\mathrm{GCN}\big(\mathbf{H}^{(2)},\boldsymbol{w}\big),$
$\mathbf{H}^{(4)}$ $=\mathrm{Linear}\big(\mathbf{H}^{(3)}\big).$
Here, $\mathbf{H}^{(\ell)}$ are node features at stage $\ell$, $\mathrm{GAT}(\cdot,\boldsymbol{w})$ and $\mathrm{GCN}(\cdot,\boldsymbol{w})$ are graph attention and graph convolution operators with edge features $\boldsymbol{w}$, $\mathrm{BN}(\cdot)$ is batch normalization, $\mathrm{ELU}(\cdot)$ is the exponential linear unit, and $\mathrm{Linear}(\cdot)$ is a trainable affine layer applied node-wise [41].
In this formulation, the first GAT layer uses the physics-guided edge weights $\boldsymbol{w}_{ij}$ to modulate directional message passing, providing a learnable counterpart to the anisotropic influence encoded by the near-field MoM interactions. The subsequent GCN layers enforce isotropic neighborhood aggregation, reflecting the smoothing and continuity behavior of $\mathbf{Z}_{\mathrm{near}}$ in (10). The final linear map produces the increment $\Delta\boldsymbol{J}^{(t)}$, which is added to the previous iterate to give
@@
providing a direct network realization of the update rule in (8).
Repeating this block for $T_{\mathrm{iter}}$ iterations yields the near-field output $\boldsymbol{J}_{\mathrm{near}}=\boldsymbol{J}^{(T_{\mathrm{iter}})}$, which serves as the finest-level source input to the far-field hierarchy.
Figure 5: Schematic of the physics-driven pipeline. The predicted current $\mathbf{J}_{\mathrm{pred}}$ bifurcates into two paths: minimizing the EFIE residual $\mathcal{L}_{phys}$ for training (red), and evaluating the RCS $\sigma$ for inference (blue).
III-B Hierarchical Far-Field Fusion Module
In the implementation, the operator $\boldsymbol{\Gamma}_{k}$ in (15) is realized by an octree-aligned pooling module.
At the finest level, the near-field graph module provides the initial source descriptor
@@ III-B Hierarchical Far-Field Fusion Module
where $\boldsymbol{S}^{(k-1)}_{c}$ denotes the feature of cluster $\mathcal{C}^{(k)}_{c}$ at level k1k-1 and $\boldsymbol{\Gamma}_{k}(\cdot)$ is a permutation-invariant pooling operator (implemented in practice by combining average and extremal statistics of the points within each cluster).
This yields a compact multilevel source hierarchy $\{\boldsymbol{S}^{(k)}\}_{k=1}^{L}$ that serves as the input to the subsequent cross-scale translation module.
The global translation operator $\boldsymbol{\Phi}_{k}$ in (16)-(21) is implemented as a shallow transformer-style block. Given the level-$k$ features $\boldsymbol{S}^{(k)}$ and the unpooled coarse representation $\mathcal{U}_{k}(\tilde{\boldsymbol{S}}^{(k-1)})$, we apply a small stack of multi-head cross-attention layers, each followed by a position-wise feed-forward network [42] and residual (with normalization). In practice, this corresponds to a two-stage stack of cross-attention followed by feed-forward refinement, providing a lightweight neural realization of the multilevel translation process described in the far-field formulation.
The disaggregation operator $\boldsymbol{\Psi}_{k}$ in (22) is implemented by a KPConv-style kernel fusion module that upsamples the coarse-level far-field current and injects the translated source features. At each level $k$, a set of learnable kernels $\{\mathbf{W}_{q}^{(k)}\}_{q=1}^{N_{\text{kern}}}$ acts on $\boldsymbol{J}_{\mathrm{far}}^{(k-1)}$ through local aggregation, and the result is combined with $\tilde{\boldsymbol{S}}^{(k)}$ as
@@
where the kernels $\mathbf{W}_{q}^{(k)}$ are learned from data and play the role of resolution-aware filters on the coarse-level current [46]. Cascading this update from $k=2$ to $L$ yields the finest-level prediction $\boldsymbol{J}_{\mathrm{far}}=\boldsymbol{J}_{\mathrm{far}}^{(L)}$, which serves as the far-field counterpart in the physics loss.
III-C Physics-Driven Optimization Mechanism
The final predicted current is obtained by summing the near- and far-field components,
$\boldsymbol{J}_{\mathrm{pred}}=\boldsymbol{J}_{\mathrm{near}}+\boldsymbol{J}_{\mathrm{far}}.$
Using $\boldsymbol{J}_{\mathrm{pred}}$, as illustrated in Fig. 5, the EFIE residual $\mathcal{L}_{phys}$ is evaluated according to the physics-based loss defined in (24). For a mini-batch of $N_{s}$ incident fields, the network computes the residual for each sample and averages them to obtain the batch loss
$\mathcal{L}=\frac{1}{N_{s}}\sum_{m=1}^{N_{s}}\mathcal{L}_{phys}^{(m)},$
which drives the end-to-end optimization of all trainable parameters. In this way, U-PINet enforces the boundary condition through batch-level residual minimization, providing a direct neural realization of the residual evaluation stage in the overall EM modeling pipeline.
Given the predicted discrete current $\boldsymbol{J}_{\mathrm{pred}}$ on $\mathcal{D}$, the scattered field and the bistatic RCS are obtained by applying the radiation operator and the far-field definition in (3) and (4), respectively.
IV Experimental Setup
@@ IV-A Data Generation
IV-B Benchmark Models and Evaluation Metrics
Three representative baseline models for EM scattering modeling, including two mainstream data-driven deep learning methods and one physics-informed hybrid model, have been implemented and compared.
@@ IV-B Benchmark Models and Evaluation Metrics
IV-C Training Configuration
@@
V Experimental Results and Discussion
In this section, we present 3D object reconstruction results with the proposed U-PINet model and compare them with baselines, using the experimental setup settings detailed in the previous Section IV.
V-A Feasibility Verification
Fig. 8 visualizes the regression performance on simple targets. The scatter points for both HH and VV polarizations tightly cluster along the diagonal $\hat{y}=y$ line, demonstrating high predictive fidelity. This visual alignment is quantitatively corroborated by Table II, where U-PINet achieves an $R^{2}$ exceeding $0.97$ and an RMSE below $0.9$ dBm for all simple 33D PEC objects (from Fig. 6), confirming the model’s robustness across different polarization states.
@@ V-A Feasibility Verification
Furthermore, Fig. 9 illustrates the RCS prediction results for the SLICY target shown in Fig. 7(a) under continuously varying observation angles. In this experiment, the elevation angle $\theta$ is fixed at $30^{\circ}$, $70^{\circ}$, $120^{\circ}$, and $150^{\circ}$, while the azimuth angle $\phi$ varies continuously from $0^{\circ}$ to $180^{\circ}$. It can be observed that the predictions of U-PINet (red curves) closely match those of MLFMA-PINN (gray curves) and the reference results (blue curves). Notably, the sharp edges and corners of the SLICY geometry induce pronounced edge diffraction effects, which introduce stronger nonlinearity and angle sensitivity into the scattering responses. In such cases, purely data-driven models (ConvNet and UNet) often struggle to stably capture the underlying electromagnetic mechanisms, whereas physics-guided approaches such as U-PINet and MLFMA-PINN demonstrate superior representation and learning capabilities.
@@
These results collectively demonstrate that U-PINet effectively encodes and preserves EM priors. It exhibits both structural fidelity and predictive accuracy, laying a solid foundation for its use as a general-purpose, PINN-based framework for EM scattering modeling. This positions U-PINet not merely as a surrogate model, but as a physically-grounded emulator capable of replacing classical solvers.
V-B Accuracy and Generalization Results
V-B1 Accuracy
We first evaluate the prediction accuracy via qualitative visualization of the induced surface currents. Fig. 12 shows the predicted current distributions for (a) the assembly body and (b) the slender forebody model. Overall, the predicted currents follow the expected physical patterns, while noticeable discrepancies mainly appear near strong coupling regions and rapidly varying edge features. Such localized errors can propagate to the far field and lead to deviations in the corresponding RCS curves.
In contrast, for the three canonical targets in Fig. 13, i.e., the cube, sphere, and cone, the predicted induced currents are almost identical to the reference solutions, and the resulting RCS responses are highly accurate. These results indicate that U-PINet can produce physically consistent and accurate induced-current predictions, which further translate into reliable RCS estimation.

### v4

@@ IV-A Near-Field Graph Encoder Module
As shown in Fig. 4(a), the physics-guided edge kernel $\mathcal{G}(\cdot)$ is realized as an attention module that takes the geometric inputs and produces a near-field edge weight $w_{ij}$ for each neighboring pair $(i,j)$. For each point $\boldsymbol{p}_{i}$, we estimate its local sampling density by Gaussian kernel [44] density estimation over $\mathcal{N}(i)$, as follows:
@@
where $N=|\mathcal{N}(i)|$ and $h>0$ is the kernel bandwidth with $\mathcal{K}(\boldsymbol{x})=\exp(-\|\boldsymbol{x}\|^{2}/2)$. The inverse density is given by:
@@
where $F_{d}(\cdot)$ is a two-layer feed-forward (FF) network. Local geometric anisotropy is encoded by the difference of unit normals $\Delta\boldsymbol{n}_{ij}=\boldsymbol{n}_{i}-\boldsymbol{n}_{j}$, yielding the normal-based attention:
@@
where $F_{n}(\cdot)$ is another two-layer FF network and $\{\gamma_{ij}\}_{j\in\mathcal{N}(i)}$ forms the normal-vector attention branch.
@@
Finally, the near-field edge weight is obtained by combining the two branches, which is compactly expressed as follows:
$w_{ij}\triangleq\mathrm{MLP}\!\left(\beta_{ij}\;\oplus\;\gamma_{ij}\right),$ (30)
where $\oplus$ denotes feature concatenation and $\mathrm{MLP}(\cdot)$ is a multi-layer perceptron. This physics-guided edge attention provides a learnable realization of $\mathcal{G}(\cdot)$ in the near-field formulation, linking local density variation and geometric anisotropy to effective self- and mutual-impedance behavior. The network implementation of the near-field operator in Section II directly follows the iterative message-passing scheme in (8)-(10). Leveraging the graph topology and physics-embedded edge weights $\boldsymbol{w}_{ij}$ established in the construction phase, each iteration realizes the kernel $h_{t}(\boldsymbol{w}_{ij},\boldsymbol{J}^{(t)}(\boldsymbol{p}_{j}))$ using a lightweight stack of attention- and convolution-based graph layers.
The latter approach is sketched in Fig. 4(b). Let $\mathbf{H}^{(0)}=\{\boldsymbol{J}^{(t)}(\boldsymbol{p}_{i})\}$ denote the initial current features. In particular, the module integrates Graph Attention Networks (GAT) [45] and Graph Convolutional Networks (GCN) [46]. One update cycle is implemented as follows:
$\mathbf{H}^{(1)}$ $\triangleq\mathrm{ELU}\!\left(\mathrm{BN}\!\left(\mathrm{GAT}\big(\mathbf{H}^{(0)},\boldsymbol{w}\big)\right)\right),$ (31)
$\mathbf{H}^{(2)}$ $\triangleq\mathrm{ELU}\!\left(\mathrm{BN}\!\left(\mathrm{GCN}\big(\mathbf{H}^{(1)},\boldsymbol{w}\big)\right)\right),$
$\mathbf{H}^{(3)}$ $\triangleq\mathrm{GCN}\big(\mathbf{H}^{(2)},\boldsymbol{w}\big),$
$\mathbf{H}^{(4)}$ $\triangleq\mathrm{Linear}\big(\mathbf{H}^{(3)}\big).$
Here, $\mathbf{H}^{(\ell)}$ are node features at stage $\ell$, $\mathrm{GAT}(\cdot,\boldsymbol{w})$ and $\mathrm{GCN}(\cdot,\boldsymbol{w})$ are graph attention and graph convolution operators with edge features $\boldsymbol{w}$, $\mathrm{BN}(\cdot)$ is batch normalization, $\mathrm{ELU}(\cdot)$ is the exponential linear unit, and $\mathrm{Linear}(\cdot)$ is a trainable affine layer applied node-wise [42]. In this formulation, the first GAT layer uses the physics-guided edge weights $\boldsymbol{w}_{ij}$’s to modulate directional message passing, providing a learnable counterpart to the anisotropic influence encoded by the near-field MoM interactions. The subsequent GCN layers enforce isotropic neighborhood aggregation, reflecting the smoothing and continuity behavior of $\mathbf{Z}_{\mathrm{near}}$ in (10). The final linear map produces the increment $\Delta\boldsymbol{J}^{(t)}$, which is added to the previous iterate to give the following expression:
@@
providing a direct network realization of (8)’s update rule. Repeating this block for $T_{\mathrm{iter}}$ iterations yields the near-field output $\boldsymbol{J}_{\mathrm{near}}=\boldsymbol{J}^{(T_{\mathrm{iter}})}$, which serves as the finest-level source input to the far-field hierarchy.
Figure 5: The proposed physics-driven pipeline. The predicted current $\mathbf{J}_{\mathrm{pred}}$ bifurcates into two paths: minimization of the EFIE residual $\mathcal{L}_{\rm phys}$ for training (red) and evaluation of the RCS $\sigma(\theta,\phi)$ for inference (blue).
IV-B Hierarchical Far-Field Fusion Module
In the implementation, the operator $\boldsymbol{\Gamma}_{k}$ in (15) is realized by an octree-aligned pooling module. At the finest level, the near-field graph module provides the initial source descriptor:
@@ IV-B Hierarchical Far-Field Fusion Module
where $\boldsymbol{S}^{(k-1)}_{c}$ denotes the feature of cluster $\mathcal{C}^{(k)}_{c}$ at level k1k-1 and $\boldsymbol{\Gamma}_{k}(\cdot)$ is a permutation-invariant pooling operator. This operator is implemented, in practice, combining average and extremal statistics of the points within each cluster. This yields a compact multilevel source hierarchy $\{\boldsymbol{S}^{(k)}\}_{k=1}^{L}$ that serves as the input to the subsequent cross-scale translation module.
The global translation operator $\boldsymbol{\Phi}_{k}$ in (16)-(21) is implemented as a shallow transformer-style block. Given the level-$k$ features $\boldsymbol{S}^{(k)}$ and the unpooled coarse representation $\mathcal{U}_{k}(\tilde{\boldsymbol{S}}^{(k-1)})$, we apply a small stack of multi-head cross-attention layers, each followed by a position-wise FF network [43] and residual (with normalization). In practice, this corresponds to a two-stage stack of cross-attention followed by FF refinement, providing a lightweight neural realization of the multilevel translation process in the far-field formulation.
The disaggregation operator $\boldsymbol{\Psi}_{k}$ in (22) is implemented by a KPConv-style kernel fusion module that upsamples the coarse-level far-field current and injects the translated source features. At each level $k$, a set of learnable kernels $\{\mathbf{W}_{q}^{(k)}\}_{q=1}^{N_{\text{kern}}}$ acts on $\boldsymbol{J}_{\mathrm{far}}^{(k-1)}$ through local aggregation, and the result is combined with $\tilde{\boldsymbol{S}}^{(k)}$ yielding the formulation:
@@
where the kernels $\mathbf{W}_{q}^{(k)}$ are learned from data and play the role of resolution-aware filters on the coarse-level current [47]. Cascading this update from $k=2$ to $L$ yields the finest-level prediction $\boldsymbol{J}_{\mathrm{far}}=\boldsymbol{J}_{\mathrm{far}}^{(L)}$, which serves as the far-field counterpart in the physics-informed loss.
IV-C Physics-Driven Optimization Mechanism
The final predicted current is obtained by summing the near- and far-field components: $\boldsymbol{J}_{\mathrm{pred}}=\boldsymbol{J}_{\mathrm{near}}+\boldsymbol{J}_{\mathrm{far}}$. Using $\boldsymbol{J}_{\mathrm{pred}}$, as illustrated in Fig. 5, the EFIE residual $\mathcal{L}_{phys}$ is evaluated according to the physics-informed residual loss in (24). For a mini-batch of $N_{s}$ incident fields, the network computes the residual for each sample and averages them to obtain the batch loss as follows:
$\mathcal{L}=\frac{1}{N_{s}}\sum_{m=1}^{N_{s}}\mathcal{L}_{\rm phys}^{(m)},$
which drives the end-to-end optimization of all trainable parameters. In this way, U-PINet enforces the boundary condition through batch-level residual minimization, providing a direct neural-network-based realization of the residual evaluation stage in the overall EM reconstruction pipeline.
Finally, using the predicted discrete current $\boldsymbol{J}_{\mathrm{pred}}$ on $\mathcal{D}$, the scattered electric field and the bistatic RCS are obtained by applying the radiation operator and the far-field definition in expressions (3) and (4), respectively.
V Experimental Setup
@@ V-A Data Generation
V-B Benchmark Models and Evaluation Metrics
Three representative baseline models for EM scattering reconstruction, including two mainstream data-driven deep learning methods and one physics-informed hybrid model, have been implemented and compared.
@@ V-B Benchmark Models and Evaluation Metrics
Unless otherwise stated, all reported metrics in Section VI are computed with respect to the MLFMA-based reference solutions, i.e., by comparing each model’s predicted currents/RCS against the corresponding MLFMA ground truth under the same geometry and excitation/observation settings.
V-C Training Configuration
@@
VI Experimental Results and Discussion
In this section, we present 3D object reconstruction results with the proposed U-PINet model and compare them with baselines, using the experimental setup settings detailed in the previous Section V.
VI-A Feasibility Verification
Fig. 8 visualizes the regression performance on simple targets. The scatter points for both HH and VV polarizations tightly cluster along the diagonal $\hat{y}=y$ line, demonstrating high predictive fidelity. This visual alignment is quantitatively corroborated by Table II, where U-PINet achieves an $R^{2}$ exceeding $0.97$ and an RMSE below $0.9$ dBm for all simple 33D PEC objects (Fig. 6), confirming the model’s robustness across different polarization states.
@@ VI-A Feasibility Verification
Furthermore, Fig. 9 illustrates the RCS prediction results for the SLICY target shown in Fig. 7(a) under continuously varying observation angles. In this experiment, the elevation angle $\theta$ is fixed at $30^{\circ}$, $70^{\circ}$, $120^{\circ}$, and $150^{\circ}$, while the azimuth angle $\phi$ varies continuously from $0^{\circ}$ to $180^{\circ}$. It can be observed that the predictions of U-PINet (red curves) closely match those of MLFMA-PINN (gray curves) and the reference results (blue curves). Notably, the sharp edges and corners of the SLICY geometry induce pronounced edge diffraction effects, which introduce stronger nonlinearity and angle sensitivity into the scattering responses. In such cases, purely data-driven models (ConvNet and UNet) often struggle to stably capture the underlying EM mechanisms, whereas physics-guided approaches such as U-PINet and MLFMA-PINN demonstrate superior representation and learning capabilities.
@@
These results collectively demonstrate that U-PINet effectively encodes and preserves EM priors. It exhibits both structural fidelity and predictive accuracy, laying a solid foundation for its use as a general-purpose, PINN-based framework for EM scattering reconstruction. This positions U-PINet not merely as a surrogate model, but as a physically-grounded emulator capable of replacing classical solvers.
VI-B Accuracy and Generalization Results
@@
Figure 14: Surface current magnitude distributions computed by MLFMA (left column) and U-PINet (right column) for the SLICY target under different incident angles. Rows (a), (b), and (c) correspond to distinct elevation angles $\theta_{\text{inc}}=0^{\circ},45^{\circ},\text{and }90^{\circ}$, respectively, at the fixed azimuth angle $\phi_{\text{inc}}=0^{\circ}$.
VI-B1 Induced-Current Reconstruction Accuracy
We first assess the reconstruction accuracy at the level of the fundamental induced surface current density $\boldsymbol{J}$. This evaluation is essential because the far-field scattered response and the resulting RCS signatures are not directly regressed by the proposed network, but are instead obtained by applying the radiation operator to the reconstructed currents. Therefore, accurate reconstruction of $\boldsymbol{J}$ serves as the primary indicator of physically consistent EM scattering reconstruction and directly determines the fidelity of the downstream RCS prediction.
As a foundational validation step, we examine the network’s performance on the three canonical targets (cube, sphere, and cone) shown in Fig. 6. Fig. 12 shows a strong qualitative agreement between the reconstructed induced currents and the MLFMA solutions, with highly consistent spatial structures. This fidelity is quantitatively corroborated by negligible error metrics: the MSEs for all three cases are consistently constrained below $8.0\times 10^{-4}$ (with average $6.5\times 10^{-4}$), while the MAPEs stabilize around $1.9\%\sim 2.5\%$. These results establish a solid baseline, confirming that U-PINet can perfectly encode the fundamental scattering mechanisms on standard geometries.
Moving to complex geometries, Fig. 13 visualizes the predicted current distributions for (a) the assembly body and (b) the slender forebody model. For the former, the main difficulty stems from pronounced EM coupling among its composite components, which results in localized interaction hot-spots and rapidly varying current structures near coupled regions. Despite this complexity, it is depicted that the predicted currents preserve the dominant spatial patterns, quantitatively achieving a MSE of $2.14\times 10^{-3}$ and a MAPE of $5.36\%$. These metrics indicate that the proposed near–far hierarchical network can effectively capture both local coupling and global radiation-consistent behavior, which aligns well with the low far-field RCS error. For the slender forebody, U-PINet similarly reconstructs physically plausible current flows over complex geometric regions, including transition zones and boundary contours, where high gradients typically challenge learning-based solvers. Specifically, it is demonstrated that the proposed network maintains an MSE of $4.85\times 10^{-3}$ and a MAPE of $9.20\%$ across the testing set. While slightly higher than that of the assembly body due to the slender geometry’s sensitivity to edge diffraction, the predicted distributions remain highly consistent with the MLFMA reference.
Finally, to evaluate robustness under varying excitation conditions, we consider the SLICY target (from Fig. 7(b)) and reconstruct the induced currents under multiple incident directions $(\theta_{\mathrm{inc}},\phi_{\mathrm{inc}})$, as illustrated in Fig. 14. This experiment keeps the geometry fixed while changing the incident angle, thereby directly testing whether the model captures the illumination-dependent scattering response. Due to its sharp edges and corner-reflector structures, SLICY exhibits pronounced edge diffraction and localized scattering singularities. As shown in Fig. 14, U-PINet closely follows these incidence-driven redistributions of the current patterns and maintains strong agreement with the MLFMA reference across the tested angles. Quantitatively, the reconstruction accuracy on SLICY achieves an MSE of $3.25\times 10^{-3}$ and a MAPE of $7.82\%$. These results indicate that U-PINet learns a generalized geometry–excitation–current mapping, rather than overfitting to a single incident configuration, and thus supports consistent downstream RCS evaluation via the same physics closure.
Collectively, these results substantiate that U-PINet achieves accurate RCS prediction via 3D EM scattering reconstruction by reconstructing the physically meaningful induced currents and subsequently evaluating the scattered fields and RCS through the governing radiation formulation.